OpenStudy (anonymous):

A triangle has sides of sqrt of 2 and 3. Which could not be the length of the third side if it is a right triangle? a) sqrt 7 b) sqrt 11 c) sqrt 13

OpenStudy (agent0smith):

 Find c, using a^2 + b^2 = c^2 But there's also another possible case:  Find a, again using a^2 + b^2 = c^2 There's not a possibility of the sqrt2 being the hypotenuse, since the hypotenuse must be the longest side, and sqrt2 is smaller than 3.